2. Julie Rios has​ take-home pay of $3,580 per month and a disability insurance policy that replaces 60 percent of earnings after a​ 90-day (3-month) waiting period. She has accumulated 80 sick days at work. Julie was involved in an auto accident and was out of work for four months. How much income did she​ lose, and how much would be replaced by her disability​ policy? How else could she replace her lost​ earnings? If after four months Julie could only return to work​ half-time for an additional three months due to continuing physical​ therapy, how might she benefit from a residual benefits​ clause? How much will she receive in disability benefits for this​ period?
Answer #1

Julie's income during disability period

monthly income* disability period in months = 3,580*4 = $14320

60% of $14320 = $8592/4 = $2148

Before applying sick day benefits and disability income​benefits, Julie would lose $2148 of income.(Round to the nearest​dollar.)